calculate the mass of an alpha particle. its binding energy is 28.2 mev
 

Grade:12

2 Answers

Arun
5 years ago

Let the mass of ‘α’ particle be xu.

α’ particle contains 2 protons and 2 neutrons.

Binding energy = (2 x1.007825 u x 1 x 1.00866 u – xu)C2 = 28.2 MeV (given).

x = 4.0016 u.

Bathina.Bhagavath Geetha Krishna
5 years ago
binding energy = \Deltam*931
\Deltam=(2*1.007825+2*1.00866)-M
\Deltam=4.032-M
BE=28.2=[4.032-M]*931
M=4.032-0.0302
M=4.002units